OpenStudy (anonymous):

After 62.0 min, 33.0% of a compound has decomposed. What is the half-life of this reaction assuming first-order kinetics?

OpenStudy (frostbite):

We set up the first order reaction rate: \[\LARGE [A]_{t}=[A]_{0}e ^{-k _{r}~t}\] As we do not know the concentration A to the time t nor the concentration to the t=0 we can use the percentage to cancel the variables. Mathematically we say: \[\LARGE [A]_{t}=[A]~0.33\] Substutute: \[\LARGE [A]_{0}*0.33=[A]_{0}e ^{-k _{r}~t}\] The [A]0 cancel and we have a expression of know variables, solve for kr as the half life can be expressed as: \[\LARGE t _{1/2}=\frac{ \ln(2) }{ k _{r} }\]

OpenStudy (frostbite):

Correction: \[\LARGE [A]_{t}=[A]_{0}~0.33\]

OpenStudy (frostbite):

wait sorry most be 0.77.

OpenStudy (frostbite):

because we know that there most be 77% to the time t.

OpenStudy (chmvijay):

 using this find K then find half time